# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v3

@@ 2. $(2)$
From now on, we assume that $A(\mathbf{x})_{1}$ is $7$-regular. Furthermore, suppose $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$. Then by II.2, we have $A\left(\mathbf{x}^{\prime}\right)_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Lemma IV.3 ensures that when given $\eta\left(A(\mathbf{x})_{1}\right)$, one can uniquely recover $A(\mathbf{x})_{1}$ from $A\left(\mathbf{x}^{\prime}\right)_{1}$. Then by Lemma IV.4, we can know the approximate positions of deletions in $A(\mathbf{x})_{1}$. This further enables us to approximately determine the positions of deletions in $A(\mathbf{x})_{i}$ for all $2\leq i\leq b$. Based on the two cases in Lemma IV.4, we split our discussion into Section IV-A and Section IV-B. Our codes for correcting two $b$-burst-deletions is presented in Theorem IV.1, whose proof is divided in into Lemmas IV.5, IV.6 and IV.7.
IV-A When Case (1) in Lemma IV.4 happens
Let $A(\mathbf{x})_{1,I_{j}}$ be the $j$th run of $A(\mathbf{x})_{1}$, where $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$ (recall that $r\left(A(\mathbf{x})_{1}\right)$ is the number of runs in $A(\mathbf{x})_{1}$) and $I_{j}\subseteq[n/b]$ is the index set of $A(\mathbf{x})_{1,I_{j}}$. For simplicity, denote $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$ and $p_{r\left(A(\mathbf{x})_{1}\right)}=n/b$. Since $A(\mathbf{x})_{1}$ is $7$-regular, we have $p_{j}-p_{j-1}=\left|I_{j}\right|\leq 7\log(n/b)$. Note that $I_{j}$ depends on $\mathbf{x}$. We omit $\mathbf{x}$ for simpler notations.
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $j_{1}<j_{2}$. Then for all $2\leq i\leq b$, one of the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $\left[p_{j_{1}-1},p_{j_{1}}\right]$ (if $j_{1}=1$, then this deletion occurred in the interval $[1,p_{1}]$) and the other occurred in the interval $\left[p_{j_{2}-1},p_{j_{2}}\right]$. So $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin\left[p_{j_{1}-1},p_{j_{1}}\right]\cup\left[p_{j_{2}-1},p_{j_{2}}\right]$. Specifically, we have
$A(\mathbf{x})_{i,k}=\begin{cases}A(\mathbf{x}^{\prime})_{i,k},\text{ if }k<p_{j_{1}-1},\\ A(\mathbf{x}^{\prime})_{i,k-1},\text{ if }p_{j_{1}}<k<p_{j_{2}-1},\\ A(\mathbf{x}^{\prime})_{i,k-2},\text{ if }k>p_{j_{2}}.\end{cases}$ (8)
for all $2\leq i\leq b$.
Therefore, we need to decode $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ from $A(\mathbf{x}^{\prime})_{i}$, for all $2\leq i\leq b$. Notice that when $j\geq 2$, the interval $[p_{j-1},p_{j}]$ intersects with both $I_{j-1}$ and $I_{j}$. Our process of decoding $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ from $A(\mathbf{x}^{\prime})_{i}$ is divided in two steps: 1) recovering $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$; 2) recovering $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$.
For the first step, for all $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$, define
$f_{j}(\mathbf{x})=\left(A(\mathbf{x})_{2,p_{j}},\ldots,A(\mathbf{x})_{b,p_{j}}\right),$
which can be viewed as an integer in $\left[0,N-1\right]$, where $N=2^{b-1}$. For uniform discussion, we also let $f_{0}(\mathbf{x})=0$. Now define $L(\mathbf{x})=f_{0}(\mathbf{x})f_{1}(\mathbf{x})\cdots f_{r\left(A(\mathbf{x})_{1}\right)}(\mathbf{x})\in[0,N-1]^{r\left(A(\mathbf{x})_{1}\right)+1}$. In other word, sequence $L(\mathbf{x})$ is the concatenation of $f_{0}(\mathbf{x}),f_{1}(\mathbf{x}),\ldots,f_{r\left(A(\mathbf{x})_{1}\right)}(\mathbf{x})$. To recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$ for all $2\leq i\leq b$, it is sufficient to obtain the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$. By Equation 8, the value of $f_{j}(\mathbf{x})$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
In the following, we always assume $j_{2}-j_{1}\geq 2$, since otherwise, the two deletions in $A(\mathbf{x})_{i}$ occurred in an interval of length at most $\left|I_{j_{1}}\cup I_{j_{2}}\right|+1\leq 14\log(n/b)+1$ and this situation is covered by case (2) in Lemma IV.4. Now we have $j_{2}-1\geq j_{1}-1+2$. Then it follows that $L(\mathbf{x})$ suffered two bursts (of length $2$) of erasures. To determine the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$, we apply Lemma II.7. Let $\mathcal{E}_{1}(\cdot)$ be the function defined in Lemma II.7. Define
$f(\mathbf{x})=\mathcal{E}_{1}\left(L(\mathbf{x})\right).$
Since $r(A(\mathbf{x})_{1})\leq n/b$, the vector $f(\mathbf{x})$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(n/b+1)+6\log N+4}=2^{\log n+O_{b}(1)}$. Now the following lemma follows from above analysis and Lemma II.7.
Lemma IV.5
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $1\leq j_{1}<j_{2}\leq r\left(A(\mathbf{x})_{1}\right)$ and $j_{2}-j_{1}\geq 2$. Suppose that $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$ are known. Then given $f(\mathbf{x})$, we can recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
Upon knowing the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$, it remains to recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$. Recall that $j_{2}-j_{1}\geq 2$. This implies $p_{j_{2}-1}>p_{j_{1}}$ and thus for all $2\leq i\leq b$, the deletion in $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ does not affect $A(\mathbf{x})_{i,I_{j_{1}}}$. Therefore, we can conclude that $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{1}}}\right)$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{2}}}\right)$, for all $2\leq i\leq b$. So we can apply Lemma IV.1 to decode $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}$, respectively.
For $1\leq j\leq r\left(A(\mathbf{x})_{1}\right)$, define
$g_{j}(\mathbf{x})=\left(\mathsf{VT}\left(A(\mathbf{x})_{2,I_{j}}\right),\cdots,\mathsf{VT}\left(A(\mathbf{x})_{b,I_{j}}\right)\right).$
Since $A(\mathbf{x})_{1}$ is $7$-regular, we have $\left|I_{j}\right|\leq 7\log(n/b)$. Then it follows from Lemma IV.1 that
$\mathsf{VT}\left(A(\mathbf{x})_{i,I_{j}}\right)\in\{0,1\}^{\log\left(7\log(n/b)+1\right)}.$
So $g_{j}(\mathbf{x})$ can be viewed as an integer in $[0,N_{2}-1]$, where $N_{2}=2^{(b-1)\log\log(n)+O_{b}(1)}$. Here, $O_{b}(1)$ is a constant dependent only on $b$.
Let $Q_{1}\geq\max\{N_{2},n/b\}$ be the smallest prime. Define
$g^{(0)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}g_{j}(\mathbf{x})\pmod{2N_{2}},$ (9)
$g^{(1)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}jg_{j}(\mathbf{x})\pmod{Q_{1}}$
Lemma IV.6
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $1\leq j_{1}<j_{2}\leq r\left(A(\mathbf{x})_{1}\right)$ and $j_{2}-j_{1}\geq 2$. Suppose that $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$ are known. If the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ are known, then given $g^{(0)}(\mathbf{x})$ and $g^{(1)}(\mathbf{x})$, we can recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
Proof:
As discussed above, we have $A(\mathbf{x}^{\prime})_{i,[p_{j_{1}-1}+1,p_{j_{1}}-1]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{1}}}\right)$ and $A(\mathbf{x}^{\prime})_{i,[p_{j_{2}-1},p_{j_{2}}-2]}\in\mathcal{B}_{1}\left(A(\mathbf{x})_{i,I_{j_{2}}}\right)$, for all $2\leq i\leq b$. So by Lemma IV.1, to recover $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$ from $\mathbf{x}^{\prime}$, it is sufficient to obtain the values of $g_{j_{1}}(\mathbf{x})$ and $g_{j_{2}}(\mathbf{x})$.
Since the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ are given, the value of $g_{j}(\mathbf{x})$ is known for all $j\neq j_{1},j_{2}$. Let $\delta^{(0)}=\left(g^{(0)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}g_{j}(\mathbf{x})\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}jg_{j}(\mathbf{x})\right)\pmod{Q_{1}}$. Then it follows from Equation 9 that
$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{2}},$ (10)
$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$ (11)
Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}$ and $0\leq\delta^{(0)}<2N_{2}$, we conclude from Equation 10 that $g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})=\delta^{(0)}$. Combining this with Equation 11, we get
$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{Q_{1}},$ (12)
$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
Since $1\leq j_{1}<j_{2}\leq n/b\leq Q_{1}$, system (12) can be uniquely solved and the unique solution is
$\left(g_{j_{1}}(\mathbf{x})\pmod{Q_{1}},g_{j_{2}}(\mathbf{x})\pmod{Q_{1}}\right).$
Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}\leq Q_{1}$, we have $g_{j_{1}}(\mathbf{x})=g_{j_{1}}(\mathbf{x})\pmod{Q_{1}}$ and $g_{j_{2}}(\mathbf{x})=g_{j_{2}}(\mathbf{x})\pmod{Q_{1}}$. ∎
IV-B When Case (2) in Lemma IV.4 happens
In this case, we can find an interval $J\subseteq[n/b]$ of length at most $21\log(n/b)$ such that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting two symbols in $A(\mathbf{x})_{1,J}$. Suppose $J=[c,d]$. Then for all $2\leq i\leq b$, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $[c-1,d]$, which is of length at most $121\log n+$ 1.
Let $P=21\log\left(n/b\right)+1$ and
Let $\xi(\cdot)$ be the function defined in Lemma II.3. For each $11\leq j\leq\left\lceil n/bP\right\rceil-$ 1, define
$h_{j}\left(\mathbf{x}\right)=\left(\xi\left(A(\mathbf{x})_{2,J_{j}}\right),\xi\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi\left(A(\mathbf{x})_{b,J_{j}}\right)\right).$
Since $\left|J_{j}\right|\leq 2P=42\log\left(n/b\right)+2$ for all $j$, each $h_{j}(\mathbf{x})$ is a binary vector of length at most $7(b-1)\log(42\log\left(n/b\right)+2)+O_{b}(1)$ and so we can view $h_{j}(\mathbf{x})$ as an integer in $\left[0,N_{3}-1\right]$, where $N_{3}=2^{7(b-1)\log(42\log\left(n/b\right)+2)+O_{b}(1)}$. Let
$\begin{array}[]{l}h_{e}(\mathbf{x})=\mathop{\sum}\limits_{j\text{ even}}h_{j}(\mathbf{x})\pmod{N_{3}},\\ h_{o}(\mathbf{x})=\mathop{\sum}\limits_{j\text{ odd}}h_{j}(\mathbf{x})\pmod{N_{3}}.\end{array}$ (13)
Lemma IV.7
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting two symbols in $A(\mathbf{x})_{1,J}$, where $J\subseteq[n/b]$ is an interval of length $21\log(n/b)$ and is known to us. Given $h_{e}\left(\mathbf{x}\right)$ and $h_{o}\left(\mathbf{x}\right)$, one can uniquely recover $A(\mathbf{x})_{i}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
Proof:
We can find a $j_{0}$ such that $J\subseteq J_{j_{0}}$. Without loss of generality, we assume that $j_{0}$ is even. Then $A(\mathbf{x})_{i,J_{j}}$ and thus $h_{j}(\mathbf{x})$ is known to us for each even $j\neq j_{0}$. So it remains to recover $A(\mathbf{x})_{J_{j_{0}}}$. If $J_{j_{0}}=[c,d]$, it is clear that $A(\mathbf{x}^{\prime})_{i,[c,d-2]}=\mathcal{B}_{2}\left(A(\mathbf{x})_{i,[c,d]}\right)$, for all $2\leq i\leq b$. By Lemma II.3, it is suffices to obtain $h_{j_{0}}(\mathbf{x})$. By Equation 13, we have
$h_{j_{0}}(\mathbf{x})\equiv\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\text{ even},j\neq j_{0}}h_{j}(\mathbf{x})\right)\pmod{N_{3}}.$
Since $0\leq h_{j_{0}}(\mathbf{x})<N_{2}$, it follows that
$h_{j_{0}}(\mathbf{x})=\left(h_{e}(\mathbf{x})-\mathop{\sum}\limits_{j\text{ even},j\neq j_{0}}h_{j}(\mathbf{x})\right)\pmod{N_{3}}.$
Now the proof is completed. ∎
Now we are ready to present our main result in this section.
Theorem IV.1
Suppose $n>2b$ and $b>1$. Let $N_{1},N_{2},N_{3},Q_{1}$, $f(\mathbf{x})$, $g^{(0)}(\mathbf{x})$, $g^{(1)}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b<N_{1}$, $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{1}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{1}$ as
$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{2}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }7\text{-regular},\\ \eta(A(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $n\geq 9$, there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $6\log n+(15b-5)\log\log n+O_{b}(1)$.
Proof:
Suppose $\mathbf{x}\in\mathcal{C}_{1}$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})$. Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},n,b)$. Then $A(\mathbf{x}^{\prime})_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Since $A(\mathbf{x})_{1}$ is $7$-regular, by Lemma IV.3, we can recover $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ based on $\eta\left(A(\mathbf{x})_{1}\right)$. Now according to Lemmas IV.4, IV.5, IV.6 and IV.7, it is clear that $\mathcal{C}$ is a $2$-$b$-burst-deletion correcting code.
Let $N_{0}=2^{4\log n+10\log\log n+O(1)}$. By Lemma IV.2, there are at least $2^{n-1}$ $7$-regular sequences when $n\geq 9$. It then follows from the pigeonhole principle that there is some $a$, $b$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that
$\left|\mathcal{C}_{1}\right|\geq\frac{2^{n-1}}{N_{0}\cdot N_{1}\cdot(2N_{2})\cdot Q_{1}\cdot(N_{3})^{2}},$
which implies by definition that
$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+\log N_{1}+\log(2N_{2})+\log Q_{1}+2\log N_{3}+1$
$=5\log n+\log Q_{1}+(15b-5)\log\log n+O_{b}(1)$
$\leq 6\log n+(15b-5)\log\log n+O_{b}(1).$
The last inequality follows from the fact $n/b\leq Q_{1}<2n/b$, which is guaranteed by Lemma II.6. ∎
IV-C Another construction when $b=2$
Recall that in Section IV-A, the decoding of $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ is conducted in two steps: 1) recovering $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$; 2) recovering $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$. This is because the interval $\left[p_{j-1},p_{j}\right]$ intersects both $I_{j-1}$ and $I_{j}$. A natural question is whether we can handle deletions in the interval $\left[p_{j-1},p_{j}\right]$ in one step, instead of two steps. Next, we show that this is possible when $b=2$.
Recall that in Section IV-A, once the first row $A(\mathbf{x})_{1}$ is recovered, the two deletions in the second row $A(\mathbf{x})_{2}$ are determined in two non-overlapping intervals $\left[p_{j_{1}-1},p_{j_{1}}\right]$ and $\left[p_{j_{2}-1},p_{j_{2}}\right]$ in $[n/2]$ if we assume $p_{j_{2}-1}>p_{j_{1}}$ (otherwise, the two deletions can be corrected using the method in Section IV-B). In Lemmas 21, 22 and 23 in [42], Lu et al proved the follows result, which is a straightforward corollary of Lemma 5 and Observation 2 in [19].
Lemma IV.8
[42, Lemma 23] Let $m$, $P_{1}$ and $P_{2}$ be positive integers, where $P_{1}+P_{2}<m$. Let $P=\min\{P_{1},P_{2}\}$. There is a function $\mathcal{E}_{2}:\Sigma_{2}^{m}\rightarrow\Sigma_{2}^{2\log m+\log P+O(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{m}$, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two bits at two known non-overlapping intervals of length at most $P_{1}$ and $P_{2}$ respectively, then given $\mathcal{E}_{2}(\mathbf{x})$ and $\mathbf{y}$, one can uniqely and efficently recover $\mathbf{x}$.
Note that $p_{j_{1}}-p_{j_{1}-1},p_{j_{2}}-p_{j_{2}-1}\leq 7\log(n/2)+1$. So, let $P_{1}=P_{2}=7\log(n/2)+1$ and $m=n/2$. Then we can apply the function $\mathcal{E}_{2}$ in Lemma IV.8 to handle interval $\left[p_{j_{1}-1},p_{j_{1}}\right]$ (and $\left[p_{j_{2}-1},p_{j_{2}}\right]$) in one step, introducing only $2\log n+\log\log n+O(1)$ additional redundant bits. Now replacing the six functions $f^{(s,\ell)}$, $g^{(}0)$ and $g^{(}1)$ in the definition of $\mathcal{C}_{1}$ by only one function $\mathcal{E}_{2}$, we have the following corollary.
Corollary IV.1
Let $b=2$ and $n\geq 9$. There is a binary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $6\log n+25\log\log n+O(1)$.
V Non-binary $2$-$b$-Burst-Deletion Correcting Codes
Let $q\geq 2$. In this section, we will extend our idea in Section IV to construct $q$-ary codes capable of correcting two $b$-burst-deletions. Surprisingly, our construction in this section gives rise to a binary code which is better than that in Theorem IV.1 when $b\geq 17$.
The first step is to establish a result similar to Lemma IV.4 to approximately determine the error positions. To that end, we associate with any $q$-ary sequence a binary sequence.
For any $x\in\Sigma_{q}$, we can uniquely write it as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<\left\lceil q/2\right\rceil$. Then for a sequence $\mathbf{x}\in\Sigma_{q}^{n}$, define $u(\mathbf{x})\triangleq u_{x_{1}}\cdots u_{x_{n}}$, i.e., $u(\mathbf{x})_{i}=u_{x_{i}}$ for all $1\leq i\leq n$. Clearly, $u(\mathbf{x})\in\Sigma_{2}^{n}$. It should be noted that this decomposition of $x\in\Sigma_{q}$ into $u_{x}$ and $v_{x}$ was also applied in [23, Section V] to construct a single-burst-deletion of variable length.
Remark V.1
@@ V Non-binary $2$-$b$-Burst-Deletion Correcting Codes
We first generalize the concept of $d$-regular sequence to sequences over general alphabets.
Definition V.1
@@
Suppose first that $q$ is even. Any $x\in\Sigma_{q}$ can be uniquely represented as $x=\frac{q}{2}u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. On the other hand, when $q$ is even, we have $\frac{q}{2}u_{x}+v_{x}\in\Sigma_{q}$ for any $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. Then it follows from Lemma IV.2 that $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}_{2,n,d}\right|(q/2)^{n}=q^{n}/2\geq q^{n-1}$.
Now suppose that $q\geq 3$ is odd. In this case, previous argument does not hold. This is because when $q$ is odd, $u_{x}=1$ and $v_{x}=(q-1)/2$, we have $\left\lceil q/2\right\rceil u_{x}+(q-1)/2=q\notin\Sigma_{q}$. Fortunately, we can apply similar ideas in proofs of [19, Lemma 11] and [22, Lemma 5] to derive our lower bound on $\left|\mathcal{R}_{q,n,d}\right|$. For $m\geq 2$, define
@@
We next prove by induction on $m$ that $\left|S_{m}^{L}\right|<(0.87q)^{m}$ for all $m\geq 2$. By the inclusion-exclusion principle, it is easy to verify that $\left|S_{2}^{L}\right|=q^{2}-\left(\frac{q+1}{2}\right)^{2}=\frac{3}{4}q^{2}-\frac{1}{2}q-\frac{1}{4}<0.75q^{2}<(0.87q)^{2}$ and $\left|S_{3}^{L}\right|=q^{3}-2q\left(\frac{q+1}{2}\right)^{2}+\left(\frac{q+1}{2}^{3}\right)=\frac{5}{8}q^{3}-\frac{5}{8}q^{2}-\frac{1}{8}q+\frac{1}{8}<0.625q^{3}<(0.87q)^{3}$. Now suppose that $m\geq 4$ and the conclusion is proved for all $m^{\prime}\leq m-2$. For $\mathbf{v}\in S_{m}^{L}$, it must holds that $v_{1}v_{2}\in S_{2}^{L}$ and $v_{3}\cdots v_{m}\in S_{m-2}^{L}$. So follows that $\left|S_{m}^{L}\right|\leq\left|S_{2}^{L}\right|\left|S_{m-2}^{L}\right|<(0.87q)^{2}(0.87q)^{m-2}=(0.87q)^{m}$.
@@
It is easy to see that $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$ and thus

### v4

@@ 2. $(2)$
We briefly explain why Lemma IV.3 is helpful. Suppose $\mathbf{x}\in\Sigma_{2}^{n}$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})$. Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},n-2b,b)$. Then $A(\mathbf{x}^{\prime})_{i}\in\mathcal{B}_{2}(A(\mathbf{x})_{i})$ for all $1\leq i\leq 2$ If $A(\mathbf{x})_{1}$ is $d$-regular (where $d\geq 7$) and $\eta(A(\mathbf{x})_{1})$ is given, Lemma IV.2 ensures that we can decode $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ and Lemma IV.3 ensures that error positions in $A(\mathbf{x})_{1}$ can be approximately determined in one or two short intervals. Then by II.2, we can approximate error positions in $A(\mathbf{x})_{i}$ within one or two short intervals, for all $2\leq i\leq b$.
However, the above analysis is valid only when $\mathbf{x}$ is a binary sequence. Since we deal with binary and non-binary cases in a unified way, it is convenient to generalize the notion of regularity to general alphabets. To that end, we associate with any $q$-ary sequence a binary sequence.
Let $q\geq 2$. For any $x\in\Sigma_{q}$, we can uniquely write it as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<\left\lceil q/2\right\rceil$. Then for a sequence $\mathbf{x}\in\Sigma_{q}^{n}$, define $u(\mathbf{x})\triangleq u_{x_{1}}\cdots u_{x_{n}}$, i.e., $u(\mathbf{x})_{i}=u_{x_{i}}$ for all $1\leq i\leq n$. Clearly, $u(\mathbf{x})\in\Sigma_{2}^{n}$ and when $q=2$, we have $u(\mathbf{x})=\mathbf{x}$. It should be noted that this decomposition of $x\in\Sigma_{q}$ into $u_{x}$ and $v_{x}$ was also applied in [23, Section V] to construct a single-burst-deletion of variable length.
Remark IV.1
@@
Definition IV.2
@@
Suppose first that $q$ is even. Any $x\in\Sigma_{q}$ can be uniquely represented as $x=\frac{q}{2}u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. On the other hand, when $q$ is even, we have $\frac{q}{2}u_{x}+v_{x}\in\Sigma_{q}$ for any $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. Then it follows from Lemma IV.1 that $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}_{2,n,d}\right|(q/2)^{n}=q^{n}/2\geq q^{n-1}$.
Now suppose that $q\geq 3$ is odd. In this case, the previous argument does not hold. This is because when $q$ is odd, $u_{x}=1$ and $v_{x}=(q-1)/2$, we have $\left\lceil q/2\right\rceil u_{x}+(q-1)/2=q\notin\Sigma_{q}$. Fortunately, we can apply similar ideas in proofs of [19, Lemma 11] and [22, Lemma 5] to derive our lower bound on $\left|\mathcal{R}_{q,n,d}\right|$. For $m\geq 2$, define
@@
We next prove by induction on $m$ that $\left|S_{m}^{L}\right|<(0.87q)^{m}$ for all $m\geq 2$. By the inclusion-exclusion principle, it is easy to verify that $\left|S_{2}^{L}\right|=q^{2}-\left(\frac{q+1}{2}\right)^{2}=\frac{3}{4}q^{2}-\frac{1}{2}q-\frac{1}{4}<0.75q^{2}<(0.87q)^{2}$ and $\left|S_{3}^{L}\right|=q^{3}-2q\left(\frac{q+1}{2}\right)^{2}+\left(\frac{q+1}{2}\right)^{3}=\frac{5}{8}q^{3}-\frac{5}{8}q^{2}-\frac{1}{8}q+\frac{1}{8}<0.625q^{3}<(0.87q)^{3}$. Now suppose that $m\geq 4$ and the conclusion is proved for all $m^{\prime}\leq m-2$. For $\mathbf{v}\in S_{m}^{L}$, it must holds that $v_{1}v_{2}\in S_{2}^{L}$ and $v_{3}\cdots v_{m}\in S_{m-2}^{L}$. So it follows that $\left|S_{m}^{L}\right|\leq\left|S_{2}^{L}\right|\left|S_{m-2}^{L}\right|<(0.87q)^{2}(0.87q)^{m-2}=(0.87q)^{m}$.
@@
It is straightforward to see that $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$ and thus